A 0.50-μF and a 1.4-μF capacitor (C1 and C2, respectively) are connected in series to a 22-V battery

Part B Calculate the charge on each capasitor. Express your answers using two significant figures separated by a comma and in the unit C

Part C Calculate the potential difference across each capacitor assuming the two capacitors are in parallel. Express your answers using two significant figures separated by a comma

Part D Calculate the charge on each capasitor assuming the two capacitors are in parallel. Express your answers using two significant figures separated by a comma.
Answer #1

b) for C2

V1 = V C2 / ( C1 + C2) = 22 * 1.4 / ( 1.4 + 0.5) = 16.21 V

q1 = C1V1 = 8.1 uC

since both are in series so they share same value of charge

q2 = q1 = 8.1 uC

======

c)

since both are in parallel so they must have same potential.

V1 = V2 = 22 V

======

d)

q1 = C1 V1 = 0.5* 22 = 11 uF

q2 = C2 V2 = 1.4* 22 = 30.8 uF

======

Comment in case any doubt, will reply for sure.. Goodluck